OpenStudy (anonymous):

((1+sin)/(1-sin)) - ((1-sin)/(1+sin)) = 4tansec (working only from one side to the other)

OpenStudy (anonymous):

\[((1+\sin x)^2 - (1-sinx)^2))/(1-\sin^2x)= (4sinx)/(\cos^2x) = 4 tanxsecx\]